#6 Find the third quartile Q3 of the list of 24 sorted values shown below. 30 31 32 33 33 33 38 39 44 45 45 47 48 49 51 59 59 59 63 65 70 76 77 79 The third quartile Q3 is______
Answer #1

data: 30 31 32 33 33 33 38 39 44 45 45 47 48 49 51 59 59 59 63 65 70 76 77 79

Excel Formula : QUARTILE(( ),1) th (n +1 term 4 Excel Formula : QUARTILE(0,2) th 2 (п + 1) Q2 term Excel Formula : QUARTILE(

here n = 24
thus, Q3 = (3*(24+1)/4)th term
Q3 = 18.75th term
Q3 = 59 + 0.25*(63-59)
Q3 = 59+1
Q3 = 60

Q3=   60 [Excel function used QUARTILE(A2:A25,3) ]